From nch@roe.ac.uk Fri Dec 19 16:42:50 2008 Date: Fri, 19 Dec 2008 16:06:45 +0000 (GMT) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Mark Holliman , Nigel Hambly , Nicholas Cross , Rob Blake , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Brian Walshe , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Peredur Williams , Bob Mann , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 19/12/08 Minutes of WFAU VDFS Science Archive meeting: 19th December 2008 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, ETWS, RPB, NJC, RSC, PMW, MSH, LGR, RGM, MAR Apologies: JPE, JMS, AL, BCW DoNM: 10am, Friday January 16th 2009 in the Vista Hut Actions discharged this week: ----------------------------- ACTION: ETWS to add a delta-mag attribute to the VSA template detection table schema. Discharged by NJC ACTION: ALL to send suggested programmes to RPB for running through the new non-survey automated curation. Discharged Actions partly discharged but continuing: ----------------------------------------- ACTION: ETWS to check out availability of SDSS-DR7 split over several files/groups/volumes Continues; ETWS notes that Univ. Chicago don't have a copy of DR7 yet; JHU have been contacted to see what the situation is regarding a multi-volume database (no reply as yet) Actions carried forward from 12/12/08 meeting: ---------------------------------------------- ACTION: All to comment on the development of the VSA metadata schema as described on the relevant TWiki page - continues ACTION: NJC to check out the incremental changes in DR7 over DR5 - continues ACTION: NCH to fold the latest PI requirements into the VSA baseline schema TWiki notes - continues ACTION: ETWS to parse that latest keyword info into the VSA metadata schema files. - continues Specific points and new actions: -------------------------------- Project management: Nothing new to report this week. WFCAM & VISTA updates: NCH noted that a call for UKIDSS observers in 09A has gone out so there is an opportunity to observe if any team member is interested. The team noted that during a VISTA update from Wil Sutherland earlier this week at the UKIDSS II meeting, SV is scheduled for late Jan/early Feb. NCH noted that a communique from Dante Minniti (VVV co-PI) has been received, and there is an open invitation to attend (via teleconf!) a meeting in Santiago on Jan 13th. NJC expressed interest in sitting in on this, as well as NCH. Comments and issues arising from CASU fortnightly minutes: No new minutes this week. Networking: ETWS noted that transfer and ingest of the final chunk of 08B data is currently underway. WSA Operations: RPB noted that the final ops test of automated non-survey release for an example small deep-stack programme has been successful. The team discussed options for processing priorities over Xmas and the New Year hols and it was decided to push all automated non-survey programmes through to the DB product ingest stage in preparation for selective release in the New Year. RPB also reported that a DR5 DXS mean astrometry correction has been applied for the new synoptic DB products. MAR noted that GPS QC for DR5 is complete up to the eyeball stage. MSH suggest we perform a full backup of the ingest DB this weekend in advance of the holiday period. RSC asked about backups of the DB-driven products (deep stacks/cats and diffs etc) and RPB noted that the LTO-3 standalone drive has been detached from djoser; the suggestion now is to stick it on thoth where we have enough administrative control to install some useful Linux backup utilities. ACTION: RPB to schedule a full WSA backup this weekend. RSC asked about the NAS box file timestamp issue, and MSH noted that it's not obvious what's going on at all; RPB volunteered to take a look. Hardware: The order for the new NAS box - "pami"; see 22nd dynasty at http://www.touregypt.net/kings.htm - is now with the supplier who have indicated that delivery will be early in the New Year. MSH and NCH noted that the migration to new server room C2 is stalled somewhat, owing to worries about the fire extinguishing system and the unavailability of the correct rackmount infrastructure. This is holding up several important enhancements and new installations, so the suggestion is to crack the whip early in the New Year. Software: RSC reported: "Mostly been tidying up loose ends before the Christmas break. Fixes of minor software quirks noticed in recent runs. Refactored CU13 to run in multiple stages which makes the logging of its current activity more clear, the time to completion prediction more reliable and to prevent it from creating empty directories unnecessarily. Also, I'm refactoring the non-survey release script to improve usability and as a test bed for VSA generalisation design plans." MAR has been working on integration of LGR's new data analysis tools into the WSA web site. Survey Data Release: DR5 is on hold, awaiting eyeball QC results from the LAS (and less importantly GPS for phase-2). NCH noted that he is pondering a source merging hack for the GPS to incorporate second-epoch K observations in the way requested by survey head Phil Lucas. Non-survey Data Release: RSC and RPB noted that all 07B non-survey programmes are now open access to the world except the two where a request from the PI to hold off was received through JAC. Regarding the roll-out of the new automated survey DB generation, this will now go ahead over Xmas with delivery of the first DBs early in the New Year. Astrogrid deployment & Data Analysis services: LGR is continuing to test the integration of his new list-driven photometry tools into the WSA web site. Miscellaneous: The fall-out from the UKIDSS II meeting earlier this week was discussed. Specific requests were o the usual FITS/VOTable output formats in addition to ASCII CSV for the new list-driven photometry tools; o a prepared DB release of the calibration data; o a "footprint" service, for accurate computation of survey areas, masks and for coverage usage modes etc; o SDSS DR7 for additional coverage in the LAS equatorial and DXS VIMOS4 fields. NCH (glass half empty) expressed caution at the meeting due to resource limitations and the fact that VISTA comes on stream in the New Year, while PMW (glass half full) noted that if the UKIDSS community make enough noise about the absolute necessity of having a particular new facility in VDFS, then at least this can be used as ammo when negotiating with our beloved research council. Overall, it was felt that VDFS put on a good show and that our users are generally very happy. Early doors: Old Bell, 5pm, to commemorate NJC's birthday amongst other things. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. =============================================================